OpenStudy (anonymous):

what is (x+3)!/(x+4)! simplified?

OpenStudy (anonymous):

hmm, If I'm not mistaken, I think it's the same as the following:\[(k+1)! = k!(k+1)\] so, as for your question : \[\frac{x!(x+3)}{x!(x+4)} = \frac{x+3}{x+4}\] not sure though ^_^

OpenStudy (anonymous):

You draw a pyramid from 1,2,3....(x-1), (x-2), (x-3),...(x)...(x+1), (x+2), (x+3), (x+4). Build a pyramid like the Egyptians did.

OpenStudy (anonymous):

._. I never used that way lol

OpenStudy (anonymous):

(x+3)!/(x+4)! = (x+3)!/ /( x+4)(x+3)!= 1/(x+4)

OpenStudy (anonymous):

(k+1)!=k!(k+1) is only true for that case. All you are doing is expanding the factorial, it's not a factorial property or anything

OpenStudy (anonymous):

thank you all!

OpenStudy (anonymous):

I think I got my pyramid wrong, the x-1's should go the other way ...(x-3), (x-2), (x-1)...